A 100 lb block rest on the floor and the coefficient of sliding friction between block and floor is 0.25. A horizontal force of 40 lb acts on the block for 3s accelerates it. compute the velocity of the block at the end of 3s


1
Expert's answer
2021-11-03T10:18:31-0400

Let's first find the acceleration of the block:


"F_h-F_{fr}=ma,""F_h-\mu_sW=\dfrac{W}{g}a,""a=\dfrac{F_h-\mu_sW}{\dfrac{W}{g}},""a=\dfrac{40\ lb-0.25\times100\ lb}{\dfrac{100\ lb}{32.17\ \dfrac{ft}{s^2}}}=4.82\ \dfrac{ft}{s^2},""a=4.82\ \dfrac{ft}{s^2}\times\dfrac{0.3048\ m}{1\ ft}=1.47\ \dfrac{m}{s^2}."

Finally, we can find the velocity of the block at the end of 3s from the kinematic equation:


"v=v_0+at=0+1.47\ \dfrac{m}{s^2}\times3\ s=4.41\ \dfrac{m}{s}."